Grade 12th pass9 grade science

22. Verify by calculating that : (a) 5 moles of CO2 and 5 moles of H2O do not have the same mass. (b) 240 g of calcium and 240 g of magnesium elements have a mole ratio of 3 : 5.

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
(a) CO2 has molar mass = 44 g mol 5 moles of CO2 have molar mass = 44×5 = 220 g H O has molar mass = 18 g mol–1 5 moles of H2O have mass = 18 × 5 g = 90 g Calculate the number of ions obtained from CaCl2 when 222 g of it is dissolved in water. Ans : 1 mole of calcium chloride = 111 g ` 222 g of CaCl2 is equivalent to 2 moles of CaCl2 Since 1 formula unit CaCl2 gives 3 ions, therefore, 1 mol of CaCl2 will give 3 moles of ions. 2 moles of CaCl2 would give 3 × 2 = 6 moles of ions No. of ions = No. of moles of ions × Avogadro number = 6 × 6.022 × 1023 = 36.132 × 1023 = 3.6132 × 1024 ions 25. Cinnabar (HgS) is a prominent ore of mercury. How many grams of mercury are present in 225 g of pure HgS? Molar mass of Hg and S are 200.6 g mol–1 and 32 g mol–1 respectively. Ans : (b) Number of moles in 240 g Ca metal 240 40 Molar mass of HgS = 200.6 + 32 Number of moles in 240 g of Mg metal 24 = 240 = 10 Ratio, 6: 10 3 : 5